# SPOT: Scalable 3D Pre-training via Occupancy Prediction for Learning Transferable 3D Representations

### v3

@@ 4.3.4 SemanticKITTI Segmentation
TABLE IX: The impact of pre-training task superiority, where we employ the detection pre-training, segmentation pre-training, occupancy pre-training, respectively. We perform fine-tuning experiments on multiple datasets of both detection and segmentation tasks, using 100% pre-training data.
@@
TABLE X: Ablation study on pre-training strategies across different datasets.
@@
TABLE XI: Experiments of extending the training schedule on nuScenes for detection task.
@@
4.4 From Supervised to Semi-supervised Pre-training
In this section, considering that SPOT requires supervised information (for the purpose of generating dense occupancy) to perform the 3D pre-training task, we study SPOT’s dependence on pre-training supervision information. In order to demonstrate SPOT’s ability to scale up, we design experiments to explore the semi-supervised setting during the pre-training phase.
Specifically, we pre-train the backbone with SPOT using only 5% sequence-level labeled data and 15% sequence-level unlabeled data, where the unlabeled data are pseudo-labeled [69] by employing a naive mean-teacher approach [70]. Refer to [70] for more details in calculating the pseudo-labels. Next, the pre-trained backbone is fine-tuned on downstream tasks including nuScenes, KITTI detection tasks, and nuScenes and SemanticKITTI segmentation tasks using different baseline models.
The experimental results of semi-supervised pre-training setting are reported in Tables VVIVII, and VIII. It can be found that semi-supervised pre-training with SPOT achieves comparable downstream performance as that of fully-supervised pre-training. It consistently improves different architectures on various datasets and tasks. Also, when incorporating more unlabeled data to perform the pre-training (e.g., comparing L5%+U5% and L5%+U15%), the performance of the downstream task significantly improves. Thus, we believe that SPOT is able to generalize to label-efficient pre-training settings and further attain performance scalability on different downstream datasets and tasks such as 3D detection and segmentation tasks.
Overall, we conclude that SPOT requires a certain amount of supervised information in the pre-training dataset, but it remains compatible with unlabeled data (as observed in Table V to VIII). Furthermore, we argue that in the context of autonomous driving, where safety is paramount, the benefits of supervised or semi-supervised pre-training outweigh the challenges of labeling. From another perspective, SPOT can alleviate the model’s reliance on human annotations in downstream tasks while achieving better model performance, thereby reducing the annotation costs associated with these downstream tasks.
@@ 4.5.1 Discussions of Pre-training
Pre-training by Different Tasks. We argue that occupancy prediction is a scalable and general task for 3D representation learning. Here we conduct experiments to compare different kinds of existing task for pre-training, including detection and segmentation tasks. Pre-training is conducted on the full Waymo dataset. Besides, fine-tuning setting employs $20\%$ KITTI data, $5\%$ nuScenes(det) data, $100\%$ SemanticKITTI data, and $100\%$ nuScenes(seg) data. The results presented in Table IX reveal that relying solely on detection as a pre-training task yields minimal performance gains, particularly when significant domain discrepancies exist, e.g. Waymo to nuScenes. Similarly, segmentation alone as a pre-training task demonstrates poor performance in the downstream detection task, likely due to the absence of localization information. On the contrary, our occupancy prediction task is beneficial to achieve consistent performance improvements for various datasets and tasks.
Fine-tuning Experiments on Extending the Training Schedule. To further demonstrate that our pre-training method enhances the backbone capacity rather than simply accelerating the convergence speed of training model, we consider conducting experiments under different training schedules. We select SECOND [11], CenterPoint [12], and DSVT [68], as the baseline method, and the experimental results are shown in Table XI. It can be seen from these results that, the results of only training 30 epochs using our SPOT pre-training can exceed the results of 150 epochs of training from scratch by $2.35\%\sim 5.78\%$.
TABLE XII: Pre-training on nuScenes and fine-tuning on KITTI for detection. We fine-tune on 20% training data.
@@
TABLE XIII: Fine-tuning performance on nuScenes benchmark for detection task based on the binary occupancy pre-training. We fine-tune on 5% training data.
@@
TABLE XIV: Fine-tuning performance of employing transformer-based structure on different datasets.
@@
TABLE XV: Fine-tuning performance on Waymo benchmark (LEVEL_2 metric). We fine-tune on 3% Waymo training data. P.D.A. represents the Pre-training Data Amount.
@@
Pre-training on nuScenes Dataset. To verify that SPOT is able to pre-train on other datasets, we utilize the model that is pre-trained on Waymo to predict occupancy labels on nuScenes dataset and generate pseudo occupancy labels. Next, we pre-train SPOT from scratch on such nuScenes data, and then fine-tune on the 20% KITTI data. As shown in Table XII, SPOT achieves significant gains compared to baseline results on KITTI dataset, demonstrating the effectiveness and generalization of SPOT.
Pre-training with Binary Occupancy Labels. We conduct additional experiments using 20% sequence-level binary occupancy-based Waymo data to perform the pre-training, and employ 5% nuScenes data for downstream fine-tuning. For consistency with previous experiments, we use the widely-adopted CenterPoint detector [12] as our baseline. The results are shown in Table XIII. It can be seen that, simple binary occupancy prediction does not bring performance gains when it performs cross-domain experiments, such as Waymo to nuScenes. This is mainly due to that, the pre-training model is difficult to learn semantically-rich information of the 3D scene when only employing the binary occupancy prediction as pre-training task.
4.5.2 Ablation Studies of SPOT
Module-level Studies. We conduct ablation experiments to analyze the individual components of the proposed SPOT. For pre-training, we uniformly sample $5\%$ Waymo data and subsequently perform fine-tuning experiments on subsets of $5\%$ nuScenes (det) data, $20\%$ KITTI data, and $20\%$ ONCE dataset, using SECOND [11] as the detector. The results presented in Table X demonstrate the effectiveness of the proposed occupancy prediction task in enhancing the performance of the downstream tasks. Moreover, our proposed strategies for pre-training, including loss balancing, beam re-sampling, and dataset balancing, yield significant improvements in different datasets.
Generalizability Studies. To further verify the generalizability of our approach towards the Transformer-based network structure, we have conducted experiments on DSVT model [68]. First, we employ the encoder of DSVT model [68] and perform the pre-training process using SPOT on 20% sequence-level data from Waymo. Then, the fine-tuning experiments are conducted on the nuScenes and ONCE datasets. The results shown in Table XIV demonstrate that, for the transformer-based baseline, SPOT also achieves significant gains under different benchmarks.
4.5.3 Discussions of Downstream Tasks
TABLE XVI: Fine-tuning performance on KITTI and nuScenes (det) benchmark with 100% data, using SECOND.
@@ 4.5.3 Discussions of Downstream Tasks
TABLE XVII: Fine-tuning performance on SemanticKITTI and nuScenes (seg) benchmark with 100% data.
@@
Figure 8: Visualizing occupancy prediction on Waymo valid.
Fine-tuning Performance on Waymo Detection. We perform detailed experiments in the downstream Waymo detection task. We evaluate the results using the official Average Precision (AP) and Average Precision with Heading (APH), with a focus on the more challenging L2-LEVEL metrics. The evaluation results on the Waymo validation set are presented in Table XV. We conduct fine-tuning on $3\%$ data using the widely adopted CenterPoint detector [12]. Furthermore, we confirm the scalability of SPOT and achieve superior performance compared to training from scratch. Specifically, SPOT improves the performance of training from scratch by 4.76% and 4.69% for CenterPoint in L2 AP and L2 APH. Table XV illustrates that SPOT with only 5% sequence-level pre-training data can outperform BEV-MAE [28] and AD-PT [25] using 100% pre-training data.
Beyond the Label-efficiency Downstream Setting. We further conduct experiments on complete downstream datasets, i.e., using 100% training data from downstream tasks to conduct the fine-tuning. The results are shown in Table XVI and Table XVII. It can be found that SPOT also achieves consistent performance gains even with $100\%$ labeled data for fine-tuning, which highlights the effectiveness of SPOT.
@@ 4.6 Visualization Results
5 Conclusion and Future Work
In this paper, we introduce SPOT, a scalable and general 3D representation learning method for LiDAR point clouds. SPOT utilizes occupancy prediction as the pre-training task and narrows domain gaps between different datasets by beam re-sampling augmentation and class-balancing strategies. Besides, we conduct a thorough theoretical analysis to uncover why the proposed occupancy pre-training task obtains temporally sufficient representations. Experimentally, consistent improvement in various downstream datasets and tasks as well as scalable pre-training are observed. We believe SPOT paves the way for large-scale pre-training on LiDAR point clouds.
@@ Acknowledgements
The research was supported by the National Key R&D Program of China (Grant No. 2022ZD0160104), the Science and Technology Commission of Shanghai Municipality (Grant No. 22DZ1100102), and Shanghai Rising Star Program (Grant No. 23QD1401000).

### v4

@@ 4.5.2 Segmentation Results
TABLE XIII: The impact of pre-training task superiority, where we employ the detection pre-training, segmentation pre-training, occupancy pre-training, respectively. We perform fine-tuning experiments on multiple datasets of both detection and segmentation tasks, using 100% pre-training data.
@@
TABLE XIV: Ablation study on pre-training strategies across different datasets.
@@
TABLE XV: Experiments of extending the training schedule on nuScenes for detection task.
@@ 4.6.1 Discussions of Pre-training
Pre-training by Different Tasks. We argue that occupancy prediction is a scalable and general task for 3D representation learning. Here we conduct experiments to compare different kinds of existing task for pre-training, including detection and segmentation tasks. Pre-training is conducted on the full Waymo dataset. Besides, fine-tuning setting employs $20\%$ KITTI data, $5\%$ nuScenes(det) data, $100\%$ SemanticKITTI data, and $100\%$ nuScenes(seg) data. The results presented in Table XIII reveal that relying solely on detection as a pre-training task yields minimal performance gains, particularly when significant domain discrepancies exist, e.g. Waymo to nuScenes. Similarly, segmentation alone as a pre-training task demonstrates poor performance in the downstream detection task, likely due to the absence of localization information. On the contrary, our occupancy prediction task is beneficial to achieve consistent performance improvements for various datasets and tasks.
Fine-tuning Experiments on Extending the Training Schedule. To further demonstrate that our pre-training method enhances the backbone capacity rather than simply accelerating the convergence speed of training model, we consider conducting experiments under different training schedules. We select SECOND [11], CenterPoint [12], and DSVT [72], as the baseline method, and the experimental results are shown in Table XV. It can be seen from these results that, the results of only training 30 epochs using our SPOT pre-training can exceed the results of 150 epochs of training from scratch by $2.35\%\sim 5.78\%$.
TABLE XVI: Pre-training on nuScenes and fine-tuning on KITTI for detection. We fine-tune on 20% training data.
@@
TABLE XVII: Fine-tuning performance on nuScenes benchmark for detection task based on the binary occupancy pre-training. We fine-tune on 5% training data.
@@
TABLE XVIII: Fine-tuning performance of employing transformer-based structure on different datasets.
@@
TABLE XIX: Fine-tuning performance on Waymo benchmark (LEVEL_2 metric). We fine-tune on 3% Waymo training data. P.D.A. represents the Pre-training Data Amount.
@@
Pre-training on nuScenes Dataset. To verify that SPOT is able to pre-train on other datasets, we utilize the model that is pre-trained on Waymo to predict occupancy labels on nuScenes dataset and generate pseudo occupancy labels. Next, we pre-train SPOT from scratch on such nuScenes data, and then fine-tune on the 20% KITTI data. As shown in Table XVI, SPOT achieves significant gains compared to baseline results on KITTI dataset, demonstrating the effectiveness and generalization of SPOT.
Pre-training with Binary Occupancy Labels. In this part, we conduct additional experiments using 20% sequence-level binary occupancy-based Waymo data to perform the pre-training, and employ 5% nuScenes data for downstream fine-tuning. For consistency with previous experiments, we use the widely-adopted CenterPoint detector [12] as our baseline. The results are shown in Table XVII. It can be seen that, simple binary occupancy prediction does not bring performance gains when it performs cross-domain experiments, such as Waymo to nuScenes. This is mainly due to that, the pre-training model is difficult to learn semantically-rich information of the 3D scene when only employing the binary occupancy prediction as pre-training task. These findings highlight the importance of carefully considering and optimizing the pre-training process to achieve superior performance in the subsequent tasks.
TABLE XX: Fine-tuning performance on KITTI and nuScenes (det) benchmark with 100% data, using SECOND.
@@
TABLE XXI: Fine-tuning performance on SemanticKITTI and nuScenes (seg) benchmark with 100% data.
@@
Figure 8: Visualizing occupancy prediction on Waymo validation set.
4.6.2 Ablation Studies of SPOT
Module-level Studies. We conduct ablation experiments to analyze the individual components of the proposed SPOT. For pre-training, we uniformly sample $5\%$ Waymo data and subsequently perform fine-tuning experiments on subsets of $5\%$ nuScenes (det) data, $20\%$ KITTI data, and $20\%$ ONCE dataset, using SECOND [11] as the detector. The results presented in Table XIV demonstrate the effectiveness of the proposed occupancy prediction task in enhancing the performance of the downstream tasks. Moreover, our proposed strategies for pre-training, including loss balancing, beam re-sampling, and dataset balancing, yield significant improvements in different datasets.
Generalizability Studies. To further verify the generalizability of our approach towards the Transformer-based network structure, we have conducted experiments on DSVT model [72]. First, we employ the encoder of DSVT model [72] and perform the pre-training process using SPOT on 20% sequence-level data from Waymo. Then, the fine-tuning experiments are conducted on the nuScenes and ONCE datasets. The results shown in Table XVIII demonstrate that, for the transformer-based baseline, SPOT also achieves significant gains under different benchmarks.
4.6.3 Discussions of Downstream Tasks
Fine-tuning Performance on Waymo Detection. We perform detailed experiments in the downstream Waymo detection task. We evaluate the results using the official Average Precision (AP) and Average Precision with Heading (APH), with a focus on the more challenging L2-LEVEL metrics. The evaluation results on the Waymo validation set are presented in Table XIX. We conduct fine-tuning on $3\%$ data using the widely adopted CenterPoint detector [12]. Furthermore, we confirm the scalability of SPOT and achieve superior performance compared to training from scratch. Specifically, SPOT improves the performance of training from scratch by 4.76% and 4.69% for CenterPoint in L2 AP and L2 APH. Table XIX illustrates that SPOT with only 5% sequence-level pre-training data can outperform BEV-MAE [28] and AD-PT [25] using 100% pre-training data.
Beyond the Label-efficiency Downstream Setting. We further conduct experiments on complete downstream datasets, i.e., using 100% training data from downstream tasks to conduct the fine-tuning. The results are shown in Table XX and Table XXI. It can be found that SPOT also achieves consistent performance gains even with $100\%$ labeled data for fine-tuning, which highlights the effectiveness of SPOT.
@@ 4.7 Visualization Results
4.8 Limitations and Future Directions
While SPOT demonstrates promising results across multiple datasets and tasks, several limitations warrant discussion:
Sensor Alignment Challenges: The practical deployment of SPOT faces engineering challenges related to camera-LiDAR calibration. While modern autonomous driving systems typically require such calibration for basic operation, achieving the precision necessary for high-quality pseudo-label generation remains a significant engineering investment, particularly for organizations without established multi-modal data pipelines.
Single-Dataset Pre-training Scope: Our current framework focuses on single-dataset pre-training to establish foundational transferability. While Fig. 5 suggests that more pre-training data could yield further improvements, multi-dataset joint pre-training introduces additional complexities including dataset fusion strategies, domain mixing effects, and potential negative transfer that require systematic investigation.
5 Conclusion
In this paper, we have introduced SPOT, a scalable and general 3D representation learning method for LiDAR point clouds. SPOT utilizes occupancy prediction as the pre-training task and narrows domain gaps between different datasets by beam re-sampling augmentation and class-balancing strategies. Besides, we conduct a thorough theoretical analysis to uncover why the proposed occupancy pre-training task obtains temporally sufficient representations. Experimentally, consistent improvement in various downstream datasets and tasks as well as scalable pre-training are observed. We believe SPOT paves the way for large-scale pre-training on LiDAR point clouds.
@@ Acknowledgements
The research was supported by Shanghai Artificial Intelligence Laboratory, the National Key R&D Program of China (Grant No. 2022ZD0160104), and the Science and Technology Commission of Shanghai Municipality (Grant No. 22DZ1100102). This work was supported by a locally commissioned task from the Shanghai Municipal Government, Shanghai Rising Star Program (Grant No. 23QD1401000), and NSFC (72342023).